# A CAV-based perimeter-free regional traffic control strategy utilizing existing parking infrastructure

### v2

@@ 4.2.2 Comparison with two perimeter control algorithms
4.2.3 Random parking locations
@@
The simulation results are shown in Figure 19. It shows that except for $i=7$, random parking locations can help foster the control performance, which is in line with our expectation. As mentioned before, when $i=9$, all parking locations are distributed at the boundary of the network, and the number of vehicles passing those areas is limited. As shown in Figure 18(d), adding randomness can help overcome this drawback and make a better use of those parking locations. Similarly, when $i=5$, the issue of short remaining travel distance existing in regular parking locations can also be tackled by the existence of parking locations farther away from the center from random distributions, as shown by Figure 18(b). When $i$ is very small, i.e., $i=3$, there are not enough parking locations to hold vehicles to generate significant improvement from random parking locations. Moreover, Figure 19(c) shows that the random parking distribution does not have significant impact on the performance when $i=7$, which implies that the regular parking location pattern is already appropriate to satisfy the need of the holding strategy to produce a favorable improvement.
@@ 4.3 Results in a partial CAV environment
Previous sections assume that all vehicles in the network are CAVs so that all vehicles can be held whenever needed. Although this assumption will not be realized in the near future, the mixed environment including both CAVs and HDVs is anticipated to be the typical condition for transportation systems. Therefore, the control performance of the proposed approach is tested in such mixed environment in which only CAVs can be held. Following the previous section, 28 centroids are randomly selected to be the parking locations for CAVs. The values {20, 40, 60, 80, 100} are tested for $\alpha$, the penetration rate of CAVs. The results are shown in Figure 20. Overall, the control performance can be improved with the increase in the penetration rate. Both Figures 20(a) and 20(b) show that a system with a penetration rate of $\alpha=80$ can generate similar performance with a full CAV environment. In addition, the marginal improvement when $4080>\alpha>$ 40 is more substantial than that when $\alpha<40$.
@@ 5 Concluding Remarks
This paper proposes a simple perimeter-free regional traffic control strategy for traffic networks with CAVs. When the network becomes congested, the proposed approach temporarily holds a subset of CAVs with long remaining travel distance at nearby parking locations to reduce the traffic burden for vehicles with short travel distance. The held vehicles will re-enter the network after a certain time period holding. Compared to the traditional perimeter control methods, which are widely used for network-level traffic control, the proposed algorithm makes the use of available parking infrastructure to partially overcome the negative impact resulting from queue accumulation from the perimeter control methods and offers more robust control with respect to demand patterns. It can also be implemented without the need to define a perimeter for the region to be protected, making it more flexible.
Micro-simulation results show that the travel time of non-held vehicles can always be reduced when the proposed strategy is enacted. However, more surprisingly and importantly, some held vehicles’ travel time can also be reduced, leading to a significant improvement in the overall operational efficiency. The superiority of the proposed algorithm over two benchmark perimeter control methods has also been demonstrated. Moreover, the control performance has been demonstrated under various configurations of parking locations and capacities. This suggests that existing excess parking capacity - no matter how small or where located - can be leveraged to improve operational performance. The results also can be used by policy makers to determine if to expand parking facilities. In addition, the results from a mixed environment, including both CAVs and HDVs, provide insight for policy making to encourage HDVs to follow this temporary holding strategy.
The thresholds for both remaining travel distance and holding time are two of the essential parameters for the proposed algorithm. It is an interesting research direction to develop theoretical optimization based models for the values of both parameters. Furthermore, testing the proposed strategy in a network with heterogeneous travel demand is promising. In this scenario, the layout of parking facilities is expected to have a more significant impact on control performance.
@@ 6 Acknowledgements
This research was supported by NSF, United States Grant CMMI-1749200.

### v3

@@ 4.2.3 Comparison with two perimeter control algorithms
4.2.4 Random parking locations with infinite parking capacity
@@
The simulation results are shown in Figure 20. It shows that except for $i=7$, random parking locations can help foster the control performance, which is in line with our expectation. As mentioned before, when $i=9$, all parking locations are distributed at the boundary of the network, and the number of vehicles passing those areas is limited. As shown in Figure 19(d), adding randomness can help overcome this drawback and make a better use of those parking locations. Similarly, when $i=5$, the issue of short remaining travel distance existing in regular parking locations can also be mitigated by the existence of parking locations farther away from the center from random distributions, as shown by Figure 19(b). When $i$ is very small, i.e., $i=3$, there are not enough parking locations to hold vehicles to generate significant improvement from random parking locations. Moreover, Figure 20(c) shows that the random parking distribution does not have significant impact on the performance when $i=7$, which implies that the regular parking location pattern is already appropriate to satisfy the need of the holding strategy to produce a favorable improvement.
@@
4.2.5 Random parking locations with parking capacity limits
To examine the impact of parking capacity under random spatial distributions of parking facilities, we apply the same capacity limits considered in Section 4.2.2 to the parking configuration shown in Figure 19(c). The corresponding results are presented in Figure 21. The findings are consistent with those obtained under the perimeter-like parking distribution (see Figure 15). Even a parking capacity of 10 spaces per facility yields substantial operational improvement compared to the baseline without the holding strategy. More importantly, a capacity limit of 20 spaces produces performance nearly identical to the scenario with no parking constraint. These results demonstrate that the proposed holding strategy does not rely on unrealistic parking availability and remains effective under more general conditions, including irregular spatial distributions and limited parking capacities.
Figure 21: Impact of parking capacity on random facility distribution.
Moreover, to validate the design choice of holding longer trips rather than shorter ones—as theoretically motivated in Section 2.2—we compare the proposed mechanism with an alternative strategy that holds vehicles whose remaining travel distance is shorter than 8 links. The comparison is conducted on the same network shown in Figure 19(c), under both infinite and limited parking capacity scenarios. The limited-capacity case is particularly important, as it ensures that the number of held vehicles is comparable between the two strategies. This allows us to isolate the effect of the holding rule itself, rather than differences arising from the total number of vehicles removed from the network.
Figure 22: Comparison between holding long trips vs. holding short trips.
The results are presented in Figure 22. Under the tested setting with $\tau=6$ mins, holding vehicles with shorter remaining distances leads to worse performance than the baseline scenario without any holding strategy. This outcome is consistent with the theoretical analysis. Holding short trips increases the average remaining travel distance of vehicles circulating in the network, thereby reducing the vehicle exit rate. Furthermore, since vehicles with longer remaining distances inherently require more time to complete their trips, re-introducing held short-trip vehicles to the network after $\tau=6$ mins may not meaningfully alleviate congestion conditions. As a result, the overall control effectiveness deteriorates.
@@ 4.3 Results in a partial CAV environment
Previous sections assume that all vehicles in the network are CAVs so that all vehicles can be held whenever needed. Although this assumption will not be realized in the near future, the mixed environment including both CAVs and HDVs is anticipated to be the typical condition for transportation systems. Therefore, the control performance of the proposed approach is tested in such mixed environment in which only CAVs can be held. Following the previous section, 28 centroids are randomly selected to be the parking locations for CAVs. The values {20, 40, 60, 80, 100} are tested for $\alpha$, the penetration rate of CAVs. The results are shown in Figure 23. Overall, the control performance can be improved with the increase in the penetration rate. Both Figures 23(a) and 23(b) show that a system with a penetration rate of $\alpha=80$ can generate similar performance with a full CAV environment. In addition, the marginal improvement when $80>\alpha>40$ is more substantial than that when $\alpha<40$.
@@ 5 Concluding Remarks
This paper proposes a simple perimeter-free regional traffic control strategy for traffic networks with CAVs. When the network becomes congested, the proposed approach temporarily holds a subset of CAVs with long remaining travel distance at nearby parking locations to reduce the traffic burden for vehicles with short travel distance. The held vehicles will re-enter the network after a certain time period holding. By selectively holding long-distance trips, the strategy generates a self-reinforcing congestion mitigation mechanism. The temporary removal of these vehicles directly reduces network accumulation. Meanwhile, because the remaining vehicles tend to have shorter remaining distances, they can complete their trips more quickly, increasing the vehicle exit rate and further accelerating congestion dissipation. This dual effect enhances traffic system recovery beyond what would be achieved through holding arbitrary or short-trip vehicles.
Compared to the traditional perimeter control methods, which are widely used for network-level traffic control, the proposed algorithm makes use of available parking infrastructure to partially overcome the negative impact resulting from queue accumulation from the perimeter control methods and offers more robust control with respect to demand patterns. It can also be implemented without the need to define a perimeter for the region to be protected, making it more flexible.
Micro-simulation results show that the travel time of non-held vehicles can always be reduced when the proposed strategy is enacted. However, surprising, the travel times of some held vehicles can also be reduced, leading to a significant improvement in the overall operational efficiency. The superiority of the proposed algorithm over two benchmark perimeter control methods has also been demonstrated. Moreover, the control performance has been demonstrated under various configurations of parking locations and capacities. This suggests that existing excess parking capacity - no matter how small or where located - can be leveraged to improve operational performance. The results also can be used by policy makers to determine if to expand parking facilities. In addition, the results from a mixed environment, including both CAVs and HDVs, provide insight for policy making to encourage HDVs to follow this temporary holding strategy.
With the benefits of the proposed method demonstrated in the findings, this paper serves as a preliminary study paving the way for future refinement of this novel traffic flow management strategy. The thresholds for both remaining travel distance and holding time are two of the essential parameters for the proposed algorithm. It is an interesting research direction to develop theoretical optimization based models for the values of both parameters.
In addition, the proposed method does not explicitly consider the interaction between reserved parking for holding and regular parking demand, which could alter traffic patterns resulting from additional cruising or unmet parking demand. Another future research direction is to jointly optimize the holding strategy and parking allocation to ensure the overall control performance, taking the influence of the holding strategy on parking demand into consideration. Moreover, in practice, incentive mechanisms such as parking fee waivers, priority re-entry, or financial compensation could improve participation. Designing dynamic incentive schemes based on trip characteristics, real-time network conditions, and expected benefits such as system travel time reduction is another important direction for future research.
@@ 6 Acknowledgements
This research was supported by NSF, United States Grants CMMI-1749200 and CMMI-2554235.